Question 2: The definition of potential of a point at a position r away from a charged particle, Q is the work done to bring a test charge from infinity to that particular location. Prove that the potential can be quantified by this equation: V -kQ/r. [Hint: use the integration rule, F. dr] Calculate the potential at a distance 0.1 mm from a point charge of 50 uC.
Answer #1

potential = kQ/r

where Q = charge

r = distance of the point from the charge

V = 9*10^9*50*10^-6 / 0.1*10^-3

= 4.5*10^9 volts